# PHLP: Sole Persistent Homology for Link Prediction - Interpretable Feature Extraction

### v1

@@ III-F Predicting the Existence of the Target Link
III-G Multiangle PHLP
@@ III-H Hybrid Method
    Persistent Image Calculation: The methods described in Sections III-C, III-D, and III-E are applied to $\mathcal{N}^{k}_{u,v}$, where $I$ denotes the PI vector. An MLP $\Phi:\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ transforms the PI into a format similar to $Z$. For the hybrid method of MA-PHLP, $\mathcal{N}^{k}_{u,v}$ is replaced with multiangle subgraphs, concatenating their PI vectors.
@@
IV Experiments
This section evaluates the performance of MA-PHLP. The experiments were also conducted using only zero-dimensional homology (MA-PHLP (dim00)). We used the area under the curve (AUC) [47] as an evaluation metric. We repeated all experiments 10 times and reported the mean and standard deviation of the AUC values.
IV-A Experimental Settings
Baselines. To evaluate the effectiveness of PHLP, we compared the proposed model with five heuristic methods, four embedding-based methods, and two GNN-based models. The heuristic methods include the Adamic-Adar (AA) [3], Katz index (Katz) [48], PageRank (PR) [49], Weisfeiler-Lehman graph kernel (WLK) [50], and Weisfeiler-Lehman neural machine (WLNM) [51]. For the embedding-based methods, we applied N2V [21], spectral clustering (SPC) [52], matrix factorization (MF) [19], and LINE [22]. Moreover, SEAL [24] and WP [28] represent the GNN-based methods.
Datasets.
TABLE I: Statistics of the datasets
Dataset #Nodes #Edges Avg. node deg. Density
USAir 332 2126 12.81 3.86e-2
NS 1589 2742 3.45 2.17e-3
PB 1222 16714 27.36 2.24e-2
Yeast 2375 11693 9.85 4.15e-3
C.ele 297 2148 14.46 4.87e-2
Power 4941 6594 2.67 5.40e-4
Router 5022 6258 2.49 4.96e-4
E.coli 1805 15660 16.24 9.61e-3
In line with previous studies [24] and [28], we evaluate the performance of our MA-PHLP on the eight datasets in Table I without node attributes: USAir [53], NS [54], PB [55], Yeast [56], C. elegans (C. ele) [57], Power [57], Router [58], and E. coli [59]. The detailed statistics for each dataset are summarized in Table I.
Implementation Details. All edges in the datasets were split into training, validation, and testing datasets with proportions of $0.85$, $0.05$, and $0.1$, respectively, ensuring a fair comparison with previous studies. The max hop $M$ was set to $3$ for most datasets (Table II). However, for the E. coli dataset, it was reduced to $2$ when employing one-dimensional homology due to memory constraints. Conversely, for the Power dataset, the max hop was set to $7$ because it does not demand heavy memory and computation time. The sigmoid function was employed for the activation function of the PHLP classifier. Tables III and IV present the results of the hybrid methods using SEAL [24] and WP [28], respectively. For these experiments, a two-layer MLP was used for the MLP $\Phi$ in Step $3$ of Section III-H. We set the $k$-hops following the original methods, SEAL and WP, and the max hops $M$ of MA-PHLP were set as the $k$, except for the Power dataset. For the Power dataset, we set the $k$-hop to $1$-hop and max hop $M$ to $7$, respectively, which is discussed in detail in Section IV-D.
TABLE II: Link prediction performance measured by the AUC on Benchmark datasets (90% observed links)
@@ IV-A Experimental Settings
SEAL $97.10\pm 0.87$ $98.25\pm 0.61$ $95.07\pm 0.39$ $97.60\pm 0.33$ $89.54\pm 1.23$ $86.21\pm 2.89$ ${95.07\pm 1.63}$ $97.57\pm 0.30$
WP $\mathbf{98.20}\pm\mathbf{0.57}$ $\mathbf{99.12}\pm\mathbf{0.45}$ $\mathbf{95.42}\pm\mathbf{0.25}$ $\mathbf{98.21}\pm\mathbf{0.17}$ $\mathbf{93.30}\pm\mathbf{0.91}$ $92.11\pm 0.76$ $\mathbf{97.15}\pm\mathbf{0.29}$ $\mathbf{98.54}\pm\mathbf{0.19}$
MA-PHLP $\underline{97.10\pm 0.69}$ $\underline{98.88\pm 0.45}$ $\underline{95.10\pm 0.26}$ $\underline{97.98\pm 0.22}$ $\underline{90.33\pm 1.16}$ $\underline{93.05\pm 0.45}$ $96.30\pm 0.43$ $97.64\pm 0.20$
MA-PHLP (dim00) $97.10\pm 0.73$ $98.78\pm 0.65$ $95.06\pm 0.28$ $97.98\pm 0.23$ $89.88\pm 1.22$ $\mathbf{93.37}\pm\mathbf{0.41}$ $\underline{96.37\pm 0.43}$ $\underline{97.72\pm 0.17}$
@@ IV-B Results
Results of MA-PHLP. Table II presents the AUC scores for each model on the benchmark datasets. Bold marks the best results, and underline indicates the second-best results. The results of AA, Katz, WLK, WLNM, N2V, SPC, MF, LINE, and SEAL are copied from SEAL [24] for comparison. The MA-PHLP demonstrates high performance across most datasets, achieving competitive scores. The proposed model outperforms several baselines, falling between the SEAL and WP models in terms of the AUC score. Notably, for the Power dataset, MA-PHLP achieves the highest AUC score, indicating its effectiveness in capturing link patterns.
@@
TABLE IV: AUC scores for WALKPOOL (WP)
with and without TDA features
@@ IV-C Ablation Study
TABLE VI: AUC scores for MA-PHLP (dim00) with various $(k,l)$-angle hops
@@ V-A Analysis of the PHLP
Figs. 6 and 7 visualize concatenated PIs to illustrate how MA-PHLP (dim0) extracts topological features for LP. We let $\mathcal{Z}\subseteq\mathbb{R}^{2\times k\times r^{2}}$ be a set of vectors calculated by MA-PHLP, where $k$ is the number of angles, and $r$ denotes the PI resolution. For $(z_{1},z_{2})\in\mathcal{Z}$, $z_{1}\in\mathbb{R}^{k\times r^{2}}$ is the concatenation of PIs for all angles with a target link, and $z_{2}\in\mathbb{R}^{k\times r^{2}}$ is the concatenation for cases without a target link. We consider a function $h:\mathbb{R}^{k\times r^{2}}\rightarrow\mathbb{R}$ defined as $h(\vec{v}_{1},...,\vec{v}_{k})=\frac{1}{k}\sum_{i=1}^{k}\lVert\vec{v}_{i}\rVert_{1}$, where $\vec{v}_{i}\in\mathbb{R}^{r^{2}}$ are PIs, and $\lVert\cdot\rVert_{1}$ denotes the $L_{1}$-norm. For visualization, we transform $\mathcal{Z}$ into points in $\mathbb{R}^{2}$ using the function $G$, defined as $G(z_{1},z_{2})=(h(z_{1}),h(z_{2}))$ for each $(z_{1},z_{2})\in\mathcal{Z}$.
@@ V-B Analysis of the Power Dataset
In most LP models, including the SOTA models SEAL and WP, the Power dataset tends to have the lowest AUC scores among the datasets. In Table II, the Power dataset is at the bottom in terms of scores across models (e.g., WLK, WLNM, MF, LINE, SEAL, and WP). However, the proposed model achieves the highest AUC scores on the Power dataset among baseline models, prompting an analysis of the reasons for this performance.
In Fig. 7, for DRNL, the Power dataset exhibits horizontal lines, indicating that the values $h(z_{2})$ have a limited range of outcomes for vectors $z_{2}$ in cases without the target link; thus, the set of values $h(z_{2})$ with the same value should be spread out. This observation implies that, for numerous subgraphs the calculation of PIs yields similar outcomes despite the differences in their topological structures, posing a challenge in distinguishing between them.
To address this problem, we applied Degree DRNL, which incorporates degree information. The points in Fig. 7 are distributed without horizontal lines, leading to the highest score increase, as listed in Table V.
@@
TABLE IX: Average number of nodes in subgraphs
for the Power and Router datasets
@@
TABLE X: Comparison of models by Max hop settings
on the Power and Router datasets
@@
However, the Power dataset does not have hub nodes, and the number of nodes in the subgraph of positive links remains small. We randomly changed the center nodes $(a,b)$ for node labeling $f^{(a,b)}_{\text{degdrnl}}$ increasing the performance, as listed in Table X. This outcome highlights that setting target nodes as the center nodes may not effectively analyze the topological structure in the case of small graphs. Furthermore, the performance for the Power dataset continues to increase with increasing hops (Table X), achieving an AUC score of $95.87$, which is significantly better than that of $92.11$ for WP.

### v2

@@ III-G Predicting the Existence of the Target Link
III-H The Interpretability of the Feature Extraction
Our method enhances the interpretability of feature extraction compared to GNN-based methods. While GNNs can extract feature vectors and visualize them using dimension reduction techniques, understanding the specific reasons behind the values of these features remains challenging. This difficulty stems from the complex and uncertain nature of the training process, which involves various factors such as optimization methods, learning rates, batch sizes, loss functions, and data distributions. In contrast, our method employs a predefined feature extraction process that operates independently of any training phase, as illustrated in Figure 1. This approach not only eliminates the ambiguities associated with training but also allows us to precisely understand what each component of the extracted feature vector represents and identify the critical aspects of our method that influence these values. This level of clarity is particularly valuable for applications requiring transparent and accountable decision-making processes.
For example, the development of the Degree DRNL node labeling was enabled by the inherent interpretability of our method. Initially, we utilized the DRNL to construct filtrations but noticed that the Power dataset exhibited the lowest accuracy among our benchmark datasets. This observation led us to conduct an extensive analysis of the subgraphs and their corresponding feature vectors within the Power dataset. As depicted in Figure 5a, we identified instances where structurally distinct subgraphs produced identical feature vectors, highlighting a limitation when utilizing DRNL. Despite one subgraph being labeled as 1 and the other being labeled as 0 according to the presence and absence of a target link, our method resulted in identical feature vectors for both subgraphs. We easily identified the reasons behind these identical vectors, which was straightforward since feature extraction does not have a training process. To address this, we incorporated degree information into the node labeling, as detailed in Equation 2. We refined the labeling by adding a fractional value between 0 and 1 based on the degree. This enhancement complements the DRNL by assigning an order according to degree, thereby prioritizing simplexes during their simultaneous addition in the filtration process. Furthermore, we established a stability theorem that guarantees minor perturbations in node labelings result in bounded variations in persistence diagrams, thus enhancing the robustness and reliability of our method.
Our methodology not only facilitates a deep understanding of the feature extraction process but also allows for targeted modifications based on specific needs. This adaptability ensures that our approach can be finely tuned to enhance performance or address specific characteristics of the data set.
III-I Multiangle PHLP
@@ III-J Hybrid Method
    Persistent Image Calculation: The methods described in Sections III-C, III-D, and III-F are applied to $\mathcal{N}^{k}_{u,v}$, where $I$ denotes the PI vector. An MLP $\Phi:\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ transforms the PI into a format similar to $Z$. For the hybrid method of MA-PHLP, $\mathcal{N}^{k}_{u,v}$ is replaced with multiangle subgraphs, concatenating their PI vectors.
@@
TABLE I: Link prediction performance measured by the AUC on Benchmark datasets (90% observed links)
@@
SEAL ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}97.10\pm 0.87}$ $98.25\pm 0.61$ ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}95.07\pm 0.39}$ $97.60\pm 0.33$ $89.54\pm 1.23$ $86.21\pm 2.89$ ${95.07\pm 1.63}$ $97.57\pm 0.30$
WP $\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}98.20\pm 0.57}$ $\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}99.12\pm 0.45}$ $\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}95.42\pm 0.25}$ $\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}98.21\pm 0.17}$ $\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}93.30\pm 0.91}$ ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}92.11\pm 0.76}$ $\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}97.15\pm 0.29}$ $\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}98.54\pm 0.19}$
LGLP $97.09\pm 0.13$ $99.12\pm 0.00$ $94.70\pm 0.04$ $97.53\pm 0.13$ $88.64\pm 0.29$ $85.63\pm 0.07$ $95.51\pm 0.07$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}98.39\pm 0.08}$
MPLP $97.01\pm 0.54$ $96.17\pm 0.84$ $94.06\pm 0.58$ $94.25\pm 0.43$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}90.48\pm 0.87}$ $73.71\pm 1.08$ $91.90\pm 0.50$ $96.67\pm 0.14$
MA-PHLP ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}97.10\pm 0.69}$ ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}98.88\pm 0.45}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}95.10\pm 0.26}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}97.98\pm 0.22}$ ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}90.33\pm 1.16}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}93.05\pm 0.45}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}96.30\pm 0.43}$ $97.64\pm 0.20$
[…]
IV Experiments
This section evaluates the performance of MA-PHLP. The experiments were also conducted using only zero-dimensional homology (MA-PHLP (dim00)). We used the area under the curve (AUC) [55] as an evaluation metric. We repeated all experiments 10 times and reported the mean and standard deviation of the AUC values. The code for our implementation is available at https://github.com/AI-hew-math/MA-PHLP.
IV-A Experimental Settings
Baselines. To evaluate the effectiveness of PHLP, we compared the proposed model with five heuristic methods, four embedding-based methods, and two GNN-based models. The heuristic methods include the Adamic-Adar (AA) [3], Katz index (Katz) [56], PageRank (PR) [57], Weisfeiler-Lehman graph kernel (WLK) [58], and Weisfeiler-Lehman neural machine (WLNM) [59]. For the embedding-based methods, we applied N2V [21], spectral clustering (SPC) [60], matrix factorization (MF) [19], and LINE [22]. Moreover, SEAL [24], WP [28], LGLP [43] and MPLP [61] represent the GNN-based methods.
Datasets. In line with previous studies [24] and [28], we evaluate the performance of our MA-PHLP on the eight datasets in Table II without node attributes: USAir [62], NS [63], PB [64], Yeast [65], C. elegans (C. ele) [66], Power [66], Router [67], and E. coli [68]. The detailed statistics for each dataset are summarized in Table II.
TABLE II: Statistics of the datasets
Dataset #Nodes #Edges Avg. node deg. Density
USAir 332 2126 12.81 3.86e-2
NS 1589 2742 3.45 2.17e-3
PB 1222 16714 27.36 2.24e-2
Yeast 2375 11693 9.85 4.15e-3
C.ele 297 2148 14.46 4.87e-2
Power 4941 6594 2.67 5.40e-4
Router 5022 6258 2.49 4.96e-4
E.coli 1805 15660 16.24 9.61e-3
Implementation Details. All edges in the datasets were split into training, validation, and testing datasets with proportions of $0.85$, $0.05$, and $0.1$, respectively, ensuring a fair comparison with previous studies. The max hop $M$ was set to $3$ for most datasets (Table I). However, for the E. coli dataset, it was reduced to $2$ when employing one-dimensional homology due to memory constraints. Conversely, for the Power dataset, the max hop was set to $7$ because it does not demand heavy memory and computation time. The sigmoid function was employed for the activation function of the PHLP classifier. Tables III and IV present the results of the hybrid methods using SEAL [24] and WP [28], respectively. We choose them because SEAL is simple and powerful GNN method and WP show the hightest scores on 7 datasets. For these experiments, a two-layer MLP was used for the MLP $\Phi$ in Step $3$ of Section III-J. We set the $k$-hops following the original methods, SEAL and WP, and the max hops $M$ of MA-PHLP were set as the $k$, except for the Power dataset. For the Power dataset, we set the $k$-hop to $1$-hop and max hop $M$ to $7$, respectively, which is discussed in detail in Section IV-D.
@@ IV-B Results
Results of MA-PHLP. Table I presents the AUC scores for each model on the benchmark datasets. The top three models are colored by First, Second, Third. The results of AA, Katz, WLK, WLNM, N2V, SPC, MF, LINE, and SEAL are copied from SEAL [24] for comparison. The MA-PHLP demonstrates high performance across most datasets, achieving competitive scores. The proposed model outperforms several baselines, falling between the GNN-based models in terms of the AUC score. Notably, for the Power dataset, MA-PHLP achieves the highest AUC score, indicating its effectiveness in capturing link patterns.
@@
TABLE IV: AUC scores for WALKPOOL (WP) with and without TDA features
@@ IV-C Ablation Study
TABLE VI: AUC scores for PHLP (dim00) with various $(k,l)$-angle hops
@@ V-A Analysis of the PHLP
Figs. 7 and 8 visualize concatenated PIs to illustrate how MA-PHLP (dim0) extracts topological features for LP. We let $\mathcal{Z}\subseteq\mathbb{R}^{2\times k\times r^{2}}$ be a set of vectors calculated by MA-PHLP, where $k$ is the number of angles, and $r$ denotes the PI resolution. For $(z_{1},z_{2})\in\mathcal{Z}$, $z_{1}\in\mathbb{R}^{k\times r^{2}}$ is the concatenation of PIs for all angles with a target link, and $z_{2}\in\mathbb{R}^{k\times r^{2}}$ is the concatenation for cases without a target link. We consider a function $h:\mathbb{R}^{k\times r^{2}}\rightarrow\mathbb{R}$ defined as $h(\vec{v}_{1},...,\vec{v}_{k})=\frac{1}{k}\sum_{i=1}^{k}\lVert\vec{v}_{i}\rVert_{1}$, where $\vec{v}_{i}\in\mathbb{R}^{r^{2}}$ are PIs, and $\lVert\cdot\rVert_{1}$ denotes the $L_{1}$-norm. For visualization, we transform $\mathcal{Z}$ into points in $\mathbb{R}^{2}$ using the function $G$, defined as $G(z_{1},z_{2})=(h(z_{1}),h(z_{2}))$ for each $(z_{1},z_{2})\in\mathcal{Z}$.
@@ V-B Analysis of the Power Dataset
In most LP models, including the SOTA models SEAL and WP, the Power dataset tends to have the lowest AUC scores among the datasets. In Table I, the Power dataset is at the bottom in terms of scores across models (e.g., WLK, WLNM, MF, LINE, SEAL, and WP). However, the proposed model achieves the highest AUC scores on the Power dataset among baseline models, prompting an analysis of the reasons for this performance.
In Figure 8, for DRNL, the Power dataset exhibits horizontal lines, indicating that the values $h(z_{2})$ have a limited range of outcomes for vectors $z_{2}$ in cases without the target link; thus, the set of values $h(z_{2})$ with the same value should be spread out. This observation implies that, for numerous subgraphs the calculation of PIs yields similar outcomes despite the differences in their topological structures, posing a challenge in distinguishing between them.
To address this problem, we applied Degree DRNL, which incorporates degree information. The points in Figure 8 are distributed without horizontal lines, leading to the highest score increase, as listed in Table V.
@@
TABLE IX: Average number of nodes in subgraphs for the Power and Router datasets
@@
TABLE X: Comparison of models by Max hop settings on the Power and Router datasets
@@
However, the Power dataset does not have hub nodes, and the number of nodes in the subgraph of positive links remains small. We randomly changed the center nodes $(a,b)$ for node labeling $f^{(a,b)}_{\text{degdrnl}}(w)$ increasing the performance, as listed in Table X. This outcome highlights that setting target nodes as the center nodes may not effectively analyze the topological structure in the case of small graphs. Furthermore, the performance for the Power dataset continues to increase with increasing hops (Table X), achieving an AUC score of $95.87$, which is significantly better than that of $92.11$ for WP.
@@ VI Conclusion
Limitations and Future works Although hybrid methods consistently improve performance, they often sacrifice interpretability. Consequently, we utilize MA-PHLP to preserve interpretability. However, MA-PHLP achieves the highest scores only on the Power dataset. To address this limitation, future efforts will be dedicated to designing hybrid approaches that maintain interpretability. These developments will ensure that our enhancements not only enhance predictive accuracy but also preserve the clarity necessary for informed decision-making in LP tasks.
Acknowledgements
This work was supported by the NRF under the grant number 2021R1A2C3009648 and partially by the NRF under the grant number MSIT (RS-2023-00219980).
Appendix A
Proof of Theorem 1. Let $G^{(u,v)}=(V,E)$ be a graph with target nodes $u,v$ in $V$ and let $f^{(u,v)}$ and $g^{(u,v)}$ be two node labeling functions based on $(u,v)$ defined on $G^{(u,v)}$. For simplicity, we denote these functions as $f$ and $g$, respectively. Denote the edge-weight functions derived from $f$ and $g$ as $W(f)$ and $W(g)$, respectively. The power set $\mathbb{X}$ of the vertex set $V$ is an abstract simplicial complex. Consider the Rips filtration function $h_{f}:\mathbb{X}\rightarrow\mathbb{R}$ defined as
$h_{f}(\sigma)=\max\{W(f)(w,z)\mid\text{ edge }\{w,z\}\subseteq\sigma\}$
whenever $p\geq 1$ for $p$-simplex $\sigma\in\mathbb{X}$, and defined as $h_{f}(z)=0$ for any 00-simplex $z$. Denote $h_{f}^{-1}((-\infty,\epsilon])$ as $K_{\epsilon}$ for $\epsilon\in\mathbb{R}$. Then, the Rips filtration is obtained as $K_{\epsilon_{1}}\hookrightarrow K_{\epsilon_{2}}\hookrightarrow\cdots\hookrightarrow K_{\epsilon_{m}}=\mathbb{X}$ for $\epsilon_{1}\leq\epsilon_{2}\leq\cdots\leq\epsilon_{m}$. By the stability theorem of the persistence diagram [53], we know that $D_{B}(dgm_{p}(f),dgm_{p}(g))\leq\lVert{h_{f}}-{h_{g}}\rVert_{\infty}$.
Now, we have $\lVert h_{f}-h_{g}\rVert_{\infty}=\lvert(h_{f}-h_{g})(\sigma_{*})\rvert$ for a $p$-simplex $\sigma_{*}\in\mathbb{X}$. By the maximality of $h_{f}$ and $h_{g}$, there are edges $(w_{f},z_{f})$ and $(w_{g},z_{g})\subseteq\sigma_{*}$ such that $h_{f}(\sigma_{*})=W(f)(w_{f},z_{f})$ and $h_{g}(\sigma_{*})=W(g)(w_{g},z_{g})$. Consider an interpolation map $\mathcal{I}(t)=(1-t)W(f)+tW(g)$ for a real number $t$, where $\mathcal{I}(0)=W(f)$ and $\mathcal{I}(1)=W(g)$. Then, define a function $\mathcal{H}(t)=\mathcal{I}(t)(w_{f},z_{f})-\mathcal{I}(t)(w_{g},z_{g})$ for real number $t$. By the maximality of the functions $h_{f}$ and $h_{g}$, we have $\mathcal{H}(0)\geq 0$ and $\mathcal{H}(1)\leq 0$. Given that $\mathcal{H}$ is continuous on the closed interval $[0,1]$, the intermediate value theorem guarantees the existence of a real number $t_{0}$ within $[0,1]$ such that $\mathcal{H}(t_{0})=0$. That is, if we write $\mathcal{I}(t_{0})$ as $W_{*}$, we have $W_{*}(w_{f},z_{f})=W_{*}(w_{g},z_{g})$. Now, observe that
$\lVert h_{f}-h_{g}\rVert_{\infty}$ $=\lvert(h_{f}-h_{g})(\sigma_{*})\rvert$
$=\lvert W(f)(w_{f},z_{f})-W(g)(w_{g},z_{g})\rvert$
$\leq\lvert W(f)(w_{f},z_{f})-W_{*}(w_{f},z_{f})\rvert$
$\indent+\lvert W_{*}(w_{g},z_{g})-W(g)(w_{g},z_{g})\rvert$
$\leq\lVert W(f)-W_{*}\rVert_{\infty}+\lVert W_{*}-W(g)\rVert_{\infty}$
$=\lVert W(f)-W(g)\rVert_{\infty}$
The last equality is derived from the expression $W_{*}=(1-t_{0})W(f)+t_{0}W(g)$.
By the definition of infinity norm, there exist two vertices $w$ and $z$ such that $\lVert W(f)-W(g)\rVert_{\infty}=\lvert W(f)(w,z)-W(g)(w,z)\rvert$. Without loss of generality, we can assume $f(w)>f(z)$. Suppose $g(w)>g(z)$. Then we have $\lVert W(f)-W(g)\rVert_{\infty}=\lvert W(f)(w,z)-W(g)(w,z)\rvert=\lvert f(w)-g(w)\rvert\leq\lVert f-g\rVert_{\infty}$.
Next, suppose $g(w)\leq g(z)$. Then we have $\lVert W(f)-W(g)\rVert_{\infty}=\lvert f(w)-g(z)\rvert$. If $f(w)-g(z)>0$, then $\lvert f(w)-g(z)\rvert=f(w)-g(z)\leq f(w)-g(w)\leq\lVert f-g\rVert_{\infty}$. If $f(w)-g(z)\leq 0$, then $\lvert f(w)-g(z)\rvert=g(z)-f(w)\leq g(z)-f(z)\leq\lVert f-g\rVert_{\infty}$. Thus we have $\lVert W(f)-W(g)\rVert_{\infty}\leq\lVert f-g\rVert_{\infty}$ and conclude $D_{B}(dgm_{p}(f),dgm_{p}(g))\leq\lVert f-g\rVert_{\infty}$. ∎
Appendix B
Algorithm of PHLP. We present the training algorithms for the PHLP and MA-PHLP, as shown in Algorithms 1 and  2.
Algorithm 1 PHLP Training Algorithm
1:Input: Observed graph $G=(V,E)$;
2:Classifier $\Phi$ with He-initialized weights $\theta$;
3:Angle hop parameters $(k,l)$, maximum epochs $Epochs$, batch size $m$;
4:Output: Trained classifier $\Phi$ with optimized weights $\theta$;
5:Sample true target links $\{x_{1},x_{2},\cdots,x_{n}\}$ from observed links $E$ and label them as 1;
6:Sample false target links $\{x_{n+1},x_{n+2},\cdots,x_{2n}\}$ from $V\times V\setminus E$ and label them as 0;
7:for $i=1\cdots 2n$ do
8:    Extract $(k,l)$-angle hop subgraph $\mathcal{N}_{x_{i}}^{(k,l)}$ for target link $x_{i}$;
9:    Compute Degree DRNL and assign edge weights to graphs with and without the target link, $\mathcal{N}^{+}$ and $\mathcal{N}^{-}$;
10:    Compute persistence diagrams $D^{+}$ and $D^{-}$ for $\mathcal{N}^{+}$ and $\mathcal{N}^{-}$ using edge weights as filtration values;
11:    Vectorize $D^{+}$ and $D^{-}$ into persistence images $PI^{+}$ and $PI^{-}$;
12:    If $k\neq l$, repeat the process for $(l,k)$-angle hop subgraph and average the two vectors;
13:    Concatenate $PI^{+}$ and $PI^{-}$ to form the final vector $PI_{i}$;
14:end for
15:Form training dataset $\mathcal{X}=\{PI_{1},PI_{2},\cdots,PI_{2n}\}$ and corresponding labels $Y=\{1,\cdots,1,0,\cdots,0\}$;
16:for $e=1\cdots Epochs$ do
17:    Sample a random batch $\{pi_{1},pi_{2},\cdots,pi_{m}\}$ from $\mathcal{X}$ and corresponding labels $\{y_{1},y_{2},\cdots,y_{m}\}$;
18:    Predict labels from classifier $\Phi$: $p_{1},p_{2},\cdots,p_{m}$;
19:    Train the classifier using binary cross-entropy loss $BCE$ and update weights $\theta$: $\nabla_{\theta}(\sum\limits_{pi_{i}\in\mathcal{X}}BCE(p_{i},y_{i}))$;
20:end for
Algorithm 2 MA-PHLP Training Algorithm
1:Input: Observed graph $G=(V,E)$;
2:Max hop $H$, maximum epochs $Epochs$, batch size $m$;
3:Classifiers $\Phi=(\Phi_{1},\cdots,\Phi_{N})$ with He-initialized weights $\theta=(\theta_{1},\cdots,\theta_{N})$;
4:Trainable parameter $\alpha=(\alpha_{1},\cdots,\alpha_{N})$, where $N=|\{(k,l)\in\mathbb{Z}^{2}:0\leq l\leq k\leq H,k>0\}|$;
5:Output: Trained classifiers $\Phi$ with optimized weights $\theta$;
6:Optimized parameter $\alpha$;
7:Sample true target links $\{x_{1},x_{2},\cdots,x_{n}\}$ from observed links $E$ and label them as 1;
8:Sample false target links $\{x_{n+1},x_{n+2},\cdots,x_{2n}\}$ from $V\times V\setminus E$ and label them as 0;
9:for $i=1\cdots 2n$ do
10:    for $k=1\cdots H$ do
11:        for $l=k\cdots H$ do
12:           Extract $(k,l)$-angle hop subgraph $\mathcal{N}_{x_{i}}^{(k,l)}$ for target link $x_{i}$;
13:           Compute Degree DRNL and assign edge weights to graphs with and without the target link, $\mathcal{N}^{+}$ and $\mathcal{N}^{-}$;
14:           Compute persistence diagrams $D^{+}$ and $D^{-}$ for $\mathcal{N}^{+}$ and $\mathcal{N}^{-}$ using edge weights as filtration values;
15:           Vectorize $D^{+}$ and $D^{-}$ into persistence images $PI^{+}$ and $PI^{-}$;
16:           If $k\neq l$, repeat the same process with the $(l,k)$-angle hop subgraph and average the two vectors;
17:           Concatenate $PI^{+}$ and $PI^{-}$ to form $PI_{(k,l)}$;
18:        end for
19:    end for
20:    $PIS_{i}=\{PI_{(k,l)}:0\leq l\leq k\leq H,k>0\}=\{PI_{1},PI_{2},\cdots,PI_{N}\}$;
21:end for
22:Form training dataset $\mathcal{X}=\{PIS_{1},PIS_{2},\cdots,PIS_{2n}\}$ and corresponding labels $Y=\{1,\cdots,1,0,\cdots,0\}$;
23:for $e=1\cdots Epochs$ do
24:    Sample a random batch $\{pis_{1},pis_{2},\cdots,pis_{m}\}$ from $\mathcal{X}$ and corresponding labels $\{y_{1},y_{2},\cdots,y_{m}\}$;
25:    Predict labels from classifiers as $p_{i}=\sum_{j=1}^{N}\alpha_{j}\Phi_{j}(pi_{j})$, where $pis_{i}=(pi_{1},\cdots,pi_{N})$;
26:    Train the classifiers and update $\alpha$ using binary cross-entropy loss $BCE$: $\nabla_{\theta,\alpha}(\sum\limits_{pis_{i}\in\mathcal{X}}BCE(p_{i},y_{i}))$;
27:end for